A body is projected with initial velocity (5i+12j)m/s from origin. Gravity acts in negative y direction. Find the horizontal range.

Grade:12

3 Answers

India VK
10 years ago
Time of flight = u * u(y) /g = 2 * 12/10 = 2.4 seconds So range = u(x) * t = 5 * 2.4 = 12 m
Mostafijur Rahaman
10 years ago
u had done a mistake the formula will be 2u/g.
Shivam Chauhan
6 years ago
X = 5t Y = 12t - (0.5) 10 t^2 solve y equation for Y=0 0 = 12 t - 5t^2 0 = t (12 -5t) t= 0 (take off) t = 2.4 going back to the X equation X = 5 t X = 5 (2.4) X = 12